How many moles of Ni are there in a sample of Ni that contains 5.40×1024 atoms?

How many Ne atoms are there in 3.80 moles of Ne?

How many moles of Ca are there in 27.0 grams of Ca?

How many grams of S are there in 2.89 moles of S?
Answer #1

ANSWER:-

1 mole = 6.022×1023  atoms,

SO 5.50 X 1024/6.022×1023 = 0.9133 X 10 = 9.133 MOLES OF NICKEL

next

1 mole = 6.022×1023

3.80 x  6.022×1023 = 22.88 x 1023 atoms of Ne

next

from the formula n=w/m

where m= mass of ca = 20 g/mol

w= 27 gm

n= 27/20 = 1.35 moles of calcium

next

from the formula n=w/m

w= n x m

mass for sulfur =32

w = 2.89 x 32 =92.48 gm of sulfur